Suppose the sample space of a statistical experiment is \lbrace A, B, C \rbrace, where P(A) = 0.87 and P(B) = 0.1. Determine:
(a) P(C) =
(b) P(A^c) =

Answer by ikleyn(52864)

.
(a)  P(C) = 1 - P(A) - P(B) = 1 - 0.87 - 0.1 = 0.03.


(b)  P(A^c) = 1 - P(A) = 1 - 0.87 = 0.13.

Solved.

And in addition, so as not to get up twice, 2 x 2 = 4.